[Hallc_running] [New Logentry] RC report Nov-30

berdnik at jlab.org berdnik at jlab.org
Tue Nov 30 08:45:01 EST 2021

Logentry Text:
We are currently taking data at the Q2=6.0, W=2.40, TH_shms=13.12deg (left) setting. We have pi- 8273 events at the moment.

We had issue with SHMS dipole at 5am, interlock was reset remotely by Steve L., downtime ~1h 

Smooth data taking over swing and owl shifts, we are planning to continue with physics until beginning of the Owl Dec-1.


Run plan for Wednesday Dec 1:

-- OWL: Configuration change to Q2=3.85, W=3.07, x=0.31, low epsilon, HMS angle=35.12 deg, HMS momentum = -1.332 GeV/c, SHMS angle= 8.39deg, SHMS momentum=-6.538 GeV/c; Data taking
-- DAY: Control access   

Pending issues:

Elevator is not working

RC meeting:

Let's have a meeting today at 10:00 AM EST.
Connection information: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3953066
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211130/75828bda/attachment.html>

More information about the Hallc_running mailing list